steven throws a ball upward with a velocity of 4.9m/s from a height of 29.4 meters when will the ball hit the ground?
 

Grade:10

1 Answers

Aman gohel
6 years ago
Case 1(when the ball remains above building)u=4.9m/s,v=0,V=u-gt=>0=4.9-9.8*t=>t=0.5sTotal time=0.5*2=>1s.Case 2:(when the ball is below building)u=4.9m/s,S=ut+0.5gt^2=>29.4=4.9*t+0.5*9.8*t^2=>t=2sTherefore total time=1+2=3s(And)